# Forming Auxiliary High-confident Instance-level Loss to Promote Learning from Label Proportions

### v1

@@ Abstract
Learning from label proportions (LLP), i.e., a challenging weakly-supervised learning task, aims to train a classifier by using bags of instances and the proportions of classes within bags, rather than annotated labels for each instance. Beyond the traditional bag-level loss, the mainstream methodology of LLP is to incorporate an auxiliary instance-level loss with pseudo-labels formed by predictions. Unfortunately, we empirically observed that the pseudo-labels are are often inaccurate due to over-smoothing, especially for the scenarios with large bag sizes, hurting the classifier induction. To alleviate this problem, we suggest a novel LLP method, namely Learning from Label Proportions with Auxiliary High-confident Instance-level Loss (L2p-ahil). Specifically, we propose a dual entropy-based weight (DEW) method to adaptively measure the confidences of pseudo-labels. It simultaneously emphasizes accurate predictions at the bag level and avoids overly smoothed predictions. We then form high-confident instance-level loss with DEW, and jointly optimize it with the bag-level loss in a self-training manner. The experimental results on benchmark datasets show that L2p-ahil can surpass the existing baseline methods, and the performance gain can be more significant as the bag size increases.
@@ 1 Introduction
During the past decades, supervised learning has been proven to gain significant success in diverse real-world applications, in which one can easily collect loads of training instances with precise supervision [10, 38]. Unfortunately, in many scenarios, it is challenging to collect a sufficient number of precisely labeled training instances due to various difficulties such as high annotation costs [16, 32] and privacy protection concerns [11]. Consequently, training instances with incomplete, inexact, and inaccurate supervision are often available only [50]. The demand for learning with such weak supervision facilitates the development of Weakly-Supervised Learning (WSL)[37], ranging many specific paradigms, e.g., semi-supervised learning [18, 49], partial label learning [24, 19], and positive-unlabeled learning [8, 12], to name just a few.
@@
Recently, an emerging paradigm of WSL, known as Learning from Label Proportions (LLP) as shown in Fig.1, has gained increasing attention within the machine learning community [48, 2, 9, 4], and it has been broadly applied in a wide range of domains such as disease diagnosis [39], population census [1], and spam email filtering [30]. Technically, LLP, as its name suggests, refers to the classification problem where a certain number of training instances are grouped into many bags with the corresponding proportion of classes, rather than annotated labels for each instance. Naturally, such sparse and inexact supervision from LLP causes significant difficulties in inducing effective classifiers for predicting labels of instances.
To handle the task of LLP, the basic method is to induce the classifier by optimizing the bag-level loss, measuring the difference between the ground-truth class proportion in a bag and the predicted class proportion formed by predictions of instances from the bag [46]. As suggested by [1], optimizing the bag-level loss only can not guarantee inducing effective classifiers. Therefore, the mainstream methodology of LLP becomes to incorporate an auxiliary instance-level loss with pseudo-labels, and jointly optimize it with the bag-level loss in a self-training manner [21, 22, 25, 7]. However, in our preliminary experiments, we empirically observed that the pseudo-labels generated by this kind of methods are often inaccurate and even meaningless, especially for the scenarios with large bag sizes. As illustrated in Fig.2, our early empirical results indicated that the pseudo-labels are inaccurate and tend to be overly smoothed, quantitatively measured by entropy values, as the bag size increases. Such low-quality pseudo-labels hurt the classifier induction, resulting in performance degradation.
To solve this problem, in this paper, we propose a novel LLP method, namely Learning Label Proportions with Auxiliary High-confident Instance-level Loss (L2p-ahil). Specifically, we suggest a dual entropy-based weighting (DEW) method to adaptively evaluate the confidence of pseudo-labels from both bag and instance levels. At the bag level, it emphasizes accurate class proportion predictions measured the difference between entropy values of ground-truth proportions and predicted proportions from the class-pivoted perspective. At the instance level, it avoids overly smoothed predictions of instances, directly measured by the entropy values. We then form a high-confident instance-level loss with DEW, and jointly optimize it with the bag-level loss in a self-training manner. The experimental results on benchmark datasets demonstrate that L2p-ahil can outperform the existing baseline methods, and the performance gain can be more significant as the bag size increases.
@@
  • To solve this problem, we propose a novel LLP method named (L2p-ahil), which adaptively measures the confidences of pseudo-labels.
@@ 2 Related Work
EPRM [46] theoretically demonstrated that an instance-level classifier can be trained using bag-level proportion information. Building on this theory, DLLP [1] proposed a bag-level loss that aligns the average of all instances’ predictions within a bag with the target distribution. Many methods have been developed based on this loss. LLP-GAN [20] introduced a generator to produce fake instances, increasing the class count to $K+1$, enabling the model to classify these fake instances as the $(K+1)^{\text{th}}$ class. LLP-VAT [40] uses a consistency regularization loss to maintain uniform predictions for an instance across various data augmentations. [26, 45] introduced self-training with a contrastive learning loss to aid training. These methods incorporate self-supervised representation learning, but without instance-level supervised training, it is difficult to train a well-performing instance-level classifier.
To improve instance-level classification performance, some approaches employ the model’s predictions as pseudo-labels for self-training. SELF-LLP [22] employs the model’s predictions from different epochs as pseudo-labels for training. PLOT and ROT [21, 7] utilize pseudo-labels generated through optimal transport for training. In the study by [25], they employed an online pseudo-labeling method with regret minimization. However, due to the significant smoothing of model predictions, especially with larger bag sizes, it naturally leads to the generation of numerous incorrect pseudo-labels. In our approach, we utilize pseudo-labels for self-training and adaptively adjust the weight of each instance’s pseudo-label based on both bag-level and instance-level entropy values. This prevents unreliable pseudo-labels from compromising the model’s performance.
From another perspective, [34, 35, 31, 28, 48, 14] are based on statistical learning theory. However, these works rely on strong assumptions that are often only applicable to specific label proportions. For instance, the method in [34, 35] heavily depends on the assumption that an example can belong to multiple bags.
@@
Figure 3: (a) The plot of the model’s pipeline with the ”Weak and Strong Augmentation” strategy. When calculating the instance-level loss $\mathcal{L}_{\mathrm{i}}$, $\mathbf{\hat{y}}^{\mathrm{s}}$ is used as the samples and $\mathbf{\tilde{y}}$ as the targets. Adaptive weights are not shown in the plot. (b) The plot of the detail process of DEW. The red and green colors represent the calculation processes in the bag level and instance level, respectively. $a$ is used to denote matrix elements for simplicity. L1-N refers to L1-Normalization.
@@
In the domain of semi-supervised learning, many methods utilize a binary adaptive weighting approach for pseudo-labels, assigning a weight of one when the maximum prediction probability exceeds a specific threshold, and zero otherwise. FixMatch [36] employs a fixed, very high threshold to identify high-confidence pseudo-labels, but this strategy inadvertently filters out many meaningful pseudo-labels. To address this issue, Dash [44] and AdaMatch [3] introduced a methodology where the threshold progressively increases as training advances. Another method employs non-binary adaptive weighting.[17] adaptively adjust the weight of each class by estimating the learning difficulty of each class. Influence [33] introduces the concept of adjusting weights using an algorithm based on the influence function. SoftMatch [5] employs a truncated Gaussian function to weight samples according to their confidence levels. [13] adaptively adjust the weight of each instance based on the entropy of the pseudo-labels for each instance. In our method, we measure the confidences of pseudo-labels based on the DEW method, adaptively adjusting the weight of each instance accordingly.
@@ 3 Proposed L2p-ahil Method
In this section, we introduce the proposed LLP method named (L2p-ahil).
@@ 3.1 Overview
Overall speaking, our L2p-ahil method is denoted as $f_{\mathbf{W}}(\cdot)$, parameterized by $\mathbf{W}$, which ingests an instance $\mathbf{x}$ and outputs its prediction $\mathbf{\hat{y}}=f_{\mathbf{W}}(\mathbf{x})$. We follow the methodology of incorporating an auxiliary instance-level loss with pseudo-labels. For the instance-level loss, we adopt the commonly used ”Weak and Strong Augmentation” strategy in the semi-supervised learning domain [36] to enhance the model’s representational learning capabilities. For each instance $\mathbf{x}$, we transform it into a weakly augmented version $\mathbf{x}^{\mathrm{w}}$ and a strongly augmented version $\mathbf{x}^{\mathrm{s}}$, where $\mathbf{x}^{\mathrm{w}}$ is specifically used to form the corresponding pseudo-label and $\mathbf{x}^{\mathrm{s}}$ is used as a training sample. The predictions for them are represented by $\mathbf{\hat{y}}^{\mathrm{w}}=f_{\mathbf{W}}(\mathbf{x}^{\mathrm{w}})$ and $\mathbf{\hat{y}}^{\mathrm{s}}=f_{\mathbf{W}}(\mathbf{x}^{\mathrm{s}})$, respectively. With the predictions of all instances, we can formulate the objective of L2p-ahil with respect to $\mathbf{W}$ as follows:
@@
where $\mathcal{L}_{\rm{b}}(\mathbf{p},\mathbf{\hat{y}}^{\mathrm{w}})$ is the bag-level proportion loss; $\mathcal{L}_{\rm{i}}(\mathbf{\tilde{y}},\mathbf{\hat{y}}^{\mathrm{s}})$ is the instance-level self-training loss and $\mathbf{\tilde{y}}$ denotes the pseudo-label; $\lambda$ is coefficient parameter used to balance the two losses.
@@
We define the pseudo-labels $\mathbf{\tilde{y}}$ as the hardened version of $\mathbf{\hat{y}}^{\mathrm{w}}$. This means that for each $\mathbf{x}_{i,j}$, its pseudo label $\mathbf{\tilde{y}}_{i,j}=[\tilde{y}_{i,j,c}]_{c=1}^{C}$ has one class with a probability of $1$, while the probabilities for all other classes are set to 0. Formally, this can be expressed as:
@@ 3.2 Dual Entropy-based Weight
where $\omega_{i,j}$ represents the adaptive weight for $\mathbf{x}_{i,j}$ and here, $c=\arg\max\mathbf{\hat{y}}^{\rm{w}}_{i,j}$ denotes the class with the highest probability. We calculate $\omega^{\rm{b}}_{i,c}$ and $\omega^{\rm{i}}_{i,j}$ based on the entropy differences between the predicted distribution and the target distribution at the bag level and instance level, respectively. The detailed process of DEW is shown in Fig. 3.
@@ 3.2.1 Bag-level Adaptive Weight
At the bag level, we construct a predicted distribution that includes the prediction results for class $c$ from all instances in $\mathbf{B}_{i}$, expressed as ${\mathbf{h}^{\prime}}_{i,c}^{\mathrm{b}}=\left[\hat{y}^{\mathrm{w}}_{i,j,c}\right]_{j=1}^{M}$. To ensure that the elements of this distribution sum up to 1, we apply L1 normalization to ${\mathbf{h}^{\prime}}_{i,c}^{\mathrm{b}}$, i.e., $\mathbf{h}_{i,c}^{\mathrm{b}}={\mathbf{h}^{\prime}}_{i,c}^{\mathrm{b}}/\lVert{\mathbf{h}^{\prime}}_{i,c}^{\mathrm{b}}\rVert_{1}$.
Based on practical knowledge, we define this target distribution as $\mathbf{\tilde{h}}^{\mathrm{b}}_{i,c}$, which is composed as follows: there are $m_{i}^{c}$ instances predicted as class $c$ with the highest probability, i.e., $1/m_{i}^{c}$, while the remaining $(M-m_{i}^{c})$ instances have the lowest probability, i.e., 0. It can be expressed as follows:
$\mathbf{\tilde{h}}^{\mathrm{b}}_{i,c}$ $=\text{L1-Normalization}\big(\big[\underbrace{\vphantom{1}1,\cdots,1}_{m_{i}^{c}},\underbrace{\vphantom{1}0,\cdots,0}_{M-m_{i}^{c}}\big]\big)$ (7)
@@
The entropy of the target distribution can then be calculated as $\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{b}}_{i,c})=\log m_{i}^{c}$. As the predicted distribution approaches the target distribution, the adaptive weight tends toward 1; conversely, it tends toward 0 when the distributions diverge. Therefore, we introduce $\sigma(x;\beta)=\exp\left(-\frac{x^{2}}{\beta}\right)$ as a mapping function, which allows us to derive the bag-level adaptive weight:
$\omega^{\mathrm{b}}_{i,c}$ $=\sigma\left(\mathcal{H}(\mathbf{h}^{\mathrm{b}}_{i,c})-\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{b}}_{i,c});\beta_{\mathrm{b}}\right)$ (8)
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{h}^{\mathrm{b}}_{i,c})-\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{b}}_{i,c})\right]^{2}}{\beta_{\mathrm{b}}}\right\}$
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{h}^{\mathrm{b}}_{i,c})-\log m_{i}^{c}\right]^{2}}{\beta_{\mathrm{b}}}\right\},$

### v2

@@ Abstract
Learning from label proportions (LLP), i.e., a challenging weakly-supervised learning task, aims to train a classifier by using bags of instances and the proportions of classes within bags, rather than annotated labels for each instance. Beyond the traditional bag-level loss, the mainstream methodology of LLP is to incorporate an auxiliary instance-level loss with pseudo-labels formed by predictions. Unfortunately, we empirically observed that the pseudo-labels are are often inaccurate due to over-smoothing, especially for the scenarios with large bag sizes, hurting the classifier induction. To alleviate this problem, we suggest a novel LLP method, namely Learning from Label Proportions with Auxiliary High-confident Instance-level Loss (L2p-ahil). Specifically, we propose a dual entropy-based weight (DEW) method to adaptively measure the confidences of pseudo-labels. It simultaneously emphasizes accurate predictions at the bag level and avoids overly smoothed predictions. We then form high-confident instance-level loss with DEW, and jointly optimize it with the bag-level loss in a self-training manner. The experimental results on benchmark datasets show that L2p-ahil can surpass the existing baseline methods, and the performance gain can be more significant as the bag size increases. The implementation of our method is available at https://github.com/TianhaoMa5/LLP-AHIL.
@@ 1 Introduction
During the past decades, supervised learning has been proven to gain significant success in diverse real-world applications, in which one can easily collect loads of training instances with precise supervision [21, 22]. Unfortunately, in many scenarios, it is challenging to collect a sufficient number of precisely labeled training instances due to various difficulties such as high annotation costs [36, 34] and privacy protection concerns [12]. Consequently, training instances with incomplete, inexact, and inaccurate supervision are often available only [50]. The demand for learning with such weak supervision facilitates the development of Weakly-Supervised Learning (WSL[40], ranging many specific paradigms, e.g., semi-supervised learning [18, 49], partial label learning [27, 20], and positive-unlabeled learning [10, 13], to name just a few.
@@
Recently, an emerging paradigm of WSL, known as Learning from Label Proportions (LLP) as shown in Fig. 1, has gained increasing attention within the machine learning community [48, 2, 11, 4], and it has been broadly applied across various computer vision domains such as remote sensing image categorization [19, 8] and medical image processing [3]. Technically, LLP, as its name suggests, refers to the classification problem where a certain number of training instances are grouped into many bags with the corresponding proportion of classes, rather than annotated labels for each instance. Naturally, such sparse and inexact supervision from LLP causes significant difficulties in inducing effective classifiers for predicting labels of instances.
To handle the task of LLP, the basic method is to induce the classifier by optimizing the bag-level loss, measuring the difference between the ground-truth class proportion in a bag and the predicted class proportion formed by predictions of instances from the bag [46]. As suggested by [23, 24], optimizing the bag-level loss only can not guarantee inducing effective classifiers. Therefore, the mainstream methodology of LLP becomes to incorporate an auxiliary instance-level loss with pseudo-labels, and jointly optimize it with the bag-level loss in a self-training manner [24, 25, 28, 9]. However, in our preliminary experiments, we empirically observed that the pseudo-labels generated by this kind of methods are often inaccurate and even meaningless, especially for the scenarios with large bag sizes. As illustrated in Fig. 2, our early empirical results indicated that the pseudo-labels are inaccurate and tend to be overly smoothed, quantitatively measured by entropy values, as the bag size increases. Such low-quality pseudo-labels hurt the classifier induction, resulting in performance degradation.
To solve this problem, in this paper, we propose a novel LLP method, namely Learning from Label Proportions with Auxiliary High-confident Instance-level Loss (L2p-ahil). Specifically, we suggest a dual entropy-based weight (DEW) method to adaptively evaluate the confidence of pseudo-labels from both bag and instance levels. At the bag level, it emphasizes accurate class proportion predictions measured the difference between entropy values of ground-truth proportions and predicted proportions from the class-pivoted perspective. At the instance level, it avoids overly smoothed predictions of instances, directly measured by the entropy values. We then form a high-confident instance-level loss with DEW, and jointly optimize it with the bag-level loss in a self-training manner. The experimental results on benchmark datasets demonstrate that L2p-ahil can outperform the existing baseline methods, and the performance gain can be more significant as the bag size increases.
@@
  • To solve this problem, we propose a novel LLP method named L2p-ahil, which adaptively measures the confidences of pseudo-labels.
@@ 2 Related Work
[46] theoretically demonstrated that an instance-level classifier can be trained using bag-level proportion information. Building on this theory, DLLP [1] proposed a bag-level loss that aligns the average of all instances’ predictions within a bag with the target distribution. Many methods have been developed based on this loss. LLP-GAN [23] introduced a generator to produce fake instances, increasing the class count to $K+1$, enabling the model to classify these fake instances as the $(K+1)^{\text{th}}$ class. LLP-VAT [41] uses a consistency regularization loss to maintain uniform predictions for an instance across various data augmentations. [29, 45] introduced self-training with a contrastive learning loss to aid training. These methods incorporate self-supervised representation learning, but without instance-level supervised training, it is difficult to train a well-performing instance-level classifier.
To improve instance-level classification performance, some approaches employ the model’s predictions as pseudo-labels for self-training. SELF-LLP [25] employs the model’s predictions from different epochs as pseudo-labels for training. ROT and PLOT [9, 24] utilize pseudo-labels generated through optimal transport for training. In the study by [28], they employed an online pseudo-labeling method with regret minimization. However, due to the significant smoothing of model predictions, especially with larger bag sizes, it naturally leads to the generation of numerous incorrect pseudo-labels. In our approach, we utilize pseudo-labels for self-training and adaptively adjust the weight of each instance’s pseudo-label based on both bag-level and instance-level entropy values. This prevents unreliable pseudo-labels from compromising the model’s performance.
From another perspective, the works of [37, 38, 33, 31, 48, 15, 5] are grounded in statistical learning theory. However, these approaches have significant limitations. For instance, the method proposed in [37, 38] heavily relies on the assumption that an example can belong to multiple bags. On the other hand, [5] introduced an unbiased estimator, but their method suffered from severe overfitting due to negative empirical loss.
@@
[…]
@@
In the domain of semi-supervised learning, many methods utilize an adaptive weighting approach for pseudo-labels. [17] adaptively adjust the weight of each class by estimating the learning difficulty of each class. Influence [35] introduces the concept of adjusting weights using an algorithm based on the influence function. SoftMatch [6] employs a truncated Gaussian function to weight samples according to their confidence levels. [14] adaptively adjust the weight of each instance based on the entropy of the pseudo-labels for each instance. However, many adaptive weighting strategies are not specifically tailored for LLP. Directly applying these methods to LLP may not yield optimal performance, as demonstrated in Sec. 4. In our method, we measure the confidences of pseudo-labels based on the DEW method, adaptively adjusting the weight of each instance accordingly.
@@ 3 Proposed L2p-ahil Method
In this section, we introduce the proposed LLP method named L2p-ahil.
@@ 3.1 Overview
Overall speaking, our L2p-ahil method is denoted as $f_{\mathbf{W}}(\cdot)$, parameterized by $\mathbf{W}$, which ingests an instance $\mathbf{x}$ and outputs its prediction $\mathbf{\hat{y}}=f_{\mathbf{W}}(\mathbf{x})$. We follow the methodology of incorporating an auxiliary instance-level loss with pseudo-labels. For the instance-level loss, we adopt the commonly used “Weak and Strong Augmentation” strategy in the semi-supervised learning domain [39] to enhance the model’s representational learning capabilities. For each instance $\mathbf{x}$, we transform it into a weakly augmented version $\mathbf{x}^{\mathrm{w}}$ and a strongly augmented version $\mathbf{x}^{\mathrm{s}}$, where $\mathbf{x}^{\mathrm{w}}$ is specifically used to form the corresponding pseudo-label and $\mathbf{x}^{\mathrm{s}}$ is used as a training sample. The predictions for them are represented by $\mathbf{\hat{y}}^{\mathrm{w}}=f_{\mathbf{W}}(\mathbf{x}^{\mathrm{w}})$ and $\mathbf{\hat{y}}^{\mathrm{s}}=f_{\mathbf{W}}(\mathbf{x}^{\mathrm{s}})$, respectively. With the predictions of all instances, we can formulate the objective of L2p-ahil with respect to $\mathbf{W}$ as follows:
@@
where $\mathcal{L}_{\rm{b}}(\mathbf{p},\mathbf{\hat{y}}^{\mathrm{w}})$ is the bag-level proportion loss; $\mathcal{L}_{\rm{i}}(\mathbf{\tilde{y}},\mathbf{\hat{y}}^{\mathrm{s}})$ is the instance-level self-training loss and $\mathbf{\tilde{y}}$ denotes the pseudo-label; $\lambda$ is coefficient parameter used to balance the two losses. The overall pipeline of the proposed model is illustrated in Fig. 3(a).
@@
We define the pseudo-labels $\mathbf{\tilde{y}}$ as the hardened version of $\mathbf{\hat{y}}^{\mathrm{w}}$. This means that for each $\mathbf{x}_{i,j}$, its pseudo label $\mathbf{\tilde{y}}_{i,j}=[\tilde{y}_{i,j,c}]_{c=1}^{C}$ has one class with a probability of $1$, while the probabilities for all other classes are set to 00. Formally, this can be expressed as:
@@ 3.2 Dual Entropy-based Weight
where $\omega_{i,j}$ represents the adaptive weight for $\mathbf{x}_{i,j}$ and here, $c=\arg\max\mathbf{\hat{y}}^{\rm{w}}_{i,j}$ denotes the class with the highest probability. We calculate $\omega^{\rm{b}}_{i,c}$ and $\omega^{\rm{i}}_{i,j}$ based on the entropy differences between the predicted distribution and the reference distribution at the bag and instance level, respectively. The detailed process of DEW is shown in Fig. 3(b).
@@ 3.2.1 Bag-level Adaptive Weight
At the bag level, the predicted distribution for class $c$ is constructed by aggregating the predictions from all instances within bag $\mathbf{B}_{i}$. Specifically, we first collect the unnormalized predictions $\left[\hat{y}^{\mathrm{w}}_{i,j,c}\right]_{j=1}^{M}$ into a vector denoted as ${\mathbf{\pi}^{\prime}}_{i,c}^{\mathrm{b}}$. To obtain a valid probability distribution (i.e., summing to $1$) at the bag level, we apply L1 normalization to ${\mathbf{\pi}^{\prime}}_{i,c}^{\mathrm{b}}$, resulting in the bag-level predicted distribution $\mathbf{\pi}_{i,c}^{\mathrm{b}}={\mathbf{\pi}^{\prime}}_{i,c}^{\mathrm{b}}/\lVert{\mathbf{\pi}^{\prime}}_{i,c}^{\mathrm{b}}\rVert_{1}$.
Based on practical knowledge, we define bag-level reference distribution as $\mathbf{\tilde{\pi}}^{\mathrm{b}}_{i,c}$, which is composed as follows: there are $m_{i}^{c}$ instances predicted as class $c$ with the highest probability, i.e., $1/m_{i}^{c}$, while the remaining $(M-m_{i}^{c})$ instances have the lowest probability, i.e., 0. It can be expressed as follows:
@@
[…]
The entropy of the bag-level reference distribution can then be calculated as $\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{b}}_{i,c})=\log m_{i}^{c}$. As the predicted distribution approaches the reference distribution, the adaptive weight tends toward 1; conversely, it tends toward 0 when the distributions diverge. Therefore, we introduce $\sigma(x;\beta)=\exp\left(-\frac{x^{2}}{\beta}\right)$ as a mapping function, which allows us to derive the bag-level adaptive weight:
$\omega^{\mathrm{b}}_{i,c}$ $=\sigma\left(\mathcal{H}(\mathbf{\pi}^{\mathrm{b}}_{i,c})-\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{b}}_{i,c});\beta_{\mathrm{b}}\right)$ (8)
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{\pi}^{\mathrm{b}}_{i,c})-\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{b}}_{i,c})\right]^{2}}{\beta_{\mathrm{b}}}\right\}$
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{\pi}^{\mathrm{b}}_{i,c})-\log m_{i}^{c}\right]^{2}}{\beta_{\mathrm{b}}}\right\},$